OpenStudy (anonymous):

4 divided by n +4=1 so 4 over n+4=1

OpenStudy (anonymous):

I believe you wrote the same thing in different words!!!!!!!!

OpenStudy (anonymous):

4/(n+4)=1?

OpenStudy (anonymous):

(4/n+4)=1 4=n+4 4-4=n 0=n If I deserve medal then give me one.